# The Illusion of State in State-Space Models

### v1

@@ Appendix A Floating-Point Arithmetic
We now extend these notions and results to matrix powering.
Definition A.3 ($\mathbb{D}$-matrix power).
For a matrix $M$ over $\mathbb{D}$ and $z\in\mathbb{Z}^{+}$, $\mathbb{D}$-matrix power is defined as
$M^{z}\,\triangleq\,\mathsf{cast}_{\mathbb{D}}(\mathsf{cast}_{\mathbb{R}}(M)^{z}).$ (9)
Mereghetti & Palano (2000) showed that when the datatype $\mathbb{D}$ is $n$-bit integers, one can compute $M^{n}$ in $\mathsf{TC}^{0}$. We note that their construction also works for computing $M^{z}$ for any $z\leq n,z\in\mathbb{Z}^{+}$. Further, as they remark, their construction can, in fact, be done in uniform $\mathsf{TC}^{0}$. Specifically, we observe most of their construction involves sums and products of constantly many $n$-bit integers, which can be done in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. The only involved step is dividing a polynomial of degree (up to) $n$ by a polynomial of degree (up to) $k-1$ and returning the remainder. It turns out that this ‘polynomial division with remainder’ operation can also be performed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (see Corollary 6.5 of Hesse et al., 2002 and an explanation in Section A.1). We thus have the following extension of Mereghetti & Palano’s result:
@@
Let $k\in\mathbb{Z}^{+}$ be a fixed constant. Let $M$ be a $k\times k$ matrix over $n$-bit integers and $z\leq n,z\in\mathbb{Z}^{+}$. Then integer matrix power $M^{z}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
We extend this result to matrix powers over $\mathbb{D}$ rather than integers:
@@
Let $k\in\mathbb{Z}^{+}$ be a fixed constant. Let $M$ be a $k\times k$ matrix over a datatype $\mathbb{D}$ with entries equivalently representable as $n$-bit integers. Let $z\leq n,z\in\mathbb{Z}^{+}$. If operators $\mathsf{cast}_{\mathbb{D}}$ and $\mathsf{cast}_{\mathbb{R}}$ are in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, then $\mathbb{D}$-matrix power $M^{z}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
@@
By preconditions of the lemma, we can compute $\mathsf{cast}_{\mathbb{R}}(M)$ in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. Since the entries of $M$ are equivalent to $n$-bit integers, $\mathsf{cast}_{R}(M)$ can be viewed as a $k\times k$ integer matrix of $n$-bit integers. By Lemma A.4, we can compute $\mathsf{cast}_{R}(M)^{z}$ using an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit. Finally, by a precondition of the lemma, we can cast the result back to $\mathbb{D}$, i.e., compute $\mathsf{cast}_{\mathbb{D}}(\mathsf{cast}_{\mathbb{R}}(M)^{z})$ which equals $M^{z}$, with an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit. ∎
See 2.4
@@
The idea is to convert (by scaling up) $M$ to another float matrix all whose entries are representable as integers, apply Lemma A.5, reverse the scaling, and cast the result back to $c\log n$ precision floats.
@@
Let $e$ be the smallest exponent across all float entries of $M$ and $q=\max\{0,-e\}$. Construct a re-scaled float matrix $\tilde{M}=M2^{q}$ by adding $q$ to the exponent of every entry of $M$, using up to $c\log n$ additional bits in the exponent if necessary to keep the computation exact. Note that $e$, $q$, and $\tilde{M}$ can easily be computed exactly by an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit as they involve fixed-arity arithmetic operations. Further, by construction, $\tilde{M}$ has non-negative exponents in all its float entries. Thus, every entry of $\tilde{M}$ represents an integer.
@@
The maximum number representable by each $c\log n$ precision float in $M$ is upper bounded by $2^{n^{c}}$. Thus, the maximum number representable by each entry of $\tilde{M}$ is $2^{n^{c}}\times 2^{q}=2^{n^{c}+q}$. Let $m=n^{c}+q$. It follows that each entry $\phi$ of $\tilde{M}$ can be equivalently represented as an $m$-bit integer. Further, this integer can be computed by left-shifting the mantissa of $\phi$ by a number of bits equal to the value of the exponent of $\phi$ (which is non-negative). Finally, this left-shift, and thus the $\mathsf{cast}_{\mathbb{R}}$ operation over $m$-precision floats, can be easily computed by an $\mathsf{L}$-uniform threshold circuit of size $\mathrm{poly}(m)$. In the other direction, casting from reals to $m$-precision floats can also be easily accomplished by an $\mathsf{L}$-uniform threshold circuit of size $\mathrm{poly}(m)$.
@@
Note that $2^{q}\in[0,n^{c}]$ and hence $m\in[n^{c},2n^{c}]$. In particular, $m\geq n$. Thus $z\leq n$ (a precision) implies $z\leq m$. Observing that $\tilde{M}$ is a matrix of floats each representable as an $m$-bit integer, we now apply Lemma A.5 with $\mathbb{D}$ being ‘float’ to conclude that float matrix power $\tilde{M}^{z}$ can be computed by an $\mathsf{L}$-uniform threshold circuit of size $\mathrm{poly}(m)$. Since $m\leq 2n^{c}$, this circuit is also of size $\mathrm{poly}(n)$.
@@
Finally, to compute $M^{z}$, we first divide each entry of $\tilde{M}^{z}$ by $2^{qz}$. This can be accomplished by subtracting $qz$ from the exponent of each entry of $\tilde{M}$; again, we do this computation exactly, using up to $(c+1)\log n$ additional bits in the exponent if necessary. We then cast all entries of the resulting matrix back to $c\log n$ precision floats. All this can be done in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, finishing the proof that $M^{z}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
A.1 $\mathsf{L}$-Uniformity of Polynomial Division in $\mathsf{TC}^{0}$
@@
Hesse et al. (2002) state that polynomial division is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ in Corollary 6.5. For historical reasons, this claim is preceded by weaker claims in older papers. We briefly clarify this situation to show why the stronger claim is valid.
@@
Reif & Tate (1992) establish that polynomial division can be done in $\mathsf{P}$-uniform $\mathsf{TC}^{0}$, whereas we state our results for $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. However, the only issue preventing these results from going through in the $\mathsf{L}$-uniform case is that, at the time of publication, it was not known whether integer division and iterated integer multiplication were computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. However, Hesse (2001) later proved exactly this. Combining the two results, Theorem 3.2 from Reif & Tate (1992) goes through with $\mathsf{L}$-uniformity. Its Corollary 3.3 then allows us to conclude that integer polynomial division can be solved by $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuits because the output of integer polynomial division is an analytic function whose Taylor expansion has a finite number of terms (Reif & Tate, 1992).
Appendix B Chess Reduction
We let $\mathcal{M}$ denote the set of chess moves in (source square, target square) notation.
See 3.3
@@ Appendix B Chess Reduction
Without loss of generality, we consider the variant of $S_{5}$ where the output is true if and only if the original first element returns to the first position after the given sequence of permutations has been applied. Given an $S_{5}$ instance, we will construct an initial board state and sequence of moves such that the final chessboard state encodes the output of the $S_{5}$ problem instance.
Initial Board State.
We construct a chessboard similar to Figure 1 but with a black rook at a8 and black queens at b8 to e8.
Chess Move Sequence.
We then construct a finite function $f:S_{5}\to\mathcal{M}^{*}$ that encodes a permutation $\pi$ as a sequence of chess moves. We first factor each permutation $\pi$ to a sequence of transpositions $\tau_{1}(\pi)\cdots\tau_{m_{\pi}}(\pi)$. Each transposition $\tau$ can in turn be expressed as a sequence of chess moves analogously to Figure 1. For example, transposing items 1 and 3 can be expressed as the move sequence: (a8, a7), (a1, b1), (c8, c6), (b1, a1), (a7, c7), (a1, b1), (c6, a6), (b1, a1), (c7, c8), (a1, b1), (a6, a8), (b1, a1), which has the crucial property that it transposes a8 with c8. We denote the mapping from transpositions to chess move sequences as $f:\mathcal{T}\to\mathcal{M}^{*}$. Putting it all together, we have that
$f(\pi)=\prod_{j=1}^{m_{\pi}}f(\tau_{j}(\pi)).$
To reduce a sequence of permutations $w\in S_{5}^{*}$, we let
$f(w)=\prod_{i=1}^{n}f(w_{i}).$
Putting It All Together.
We call our oracle for chess state tracking with the constructed initial board state and $f(w)$ as the sequence of chess moves. By construction, we can then return true if and only if the rook is at a8. The reduction can be implemented in $\mathsf{NC}^{0}$ because it is a simple elementwise mapping of the input tokens, and decoding from the output chessboard is a finite table lookup. ∎
As a fun aside, we note that the chess board constructed in Proposition 3.3 is reachable in a standard chess game. The chess sequences encoding permutation sequences are all valid chess games, except that they ignore the fact that repeated board states in chess will technically lead to draws.

### v2

@@ A.1 Complexity of Iterated Scalar Multiplication
A.2 Complexity of Matrix Powering
The second special case we analyze is matrix powering: i.e., a matrix product where all the matrices being powered are the same. Mereghetti & Palano (2000) showed that when the datatype $\mathbb{D}$ is $n$-bit integers, one can compute $\mathbf{M}^{n}$ in $\mathsf{TC}^{0}$. We note that their construction also works for computing $\mathbf{M}^{z}$ for any $z\leq n,z\in\mathbb{Z}^{+}$. Further, as they remark, their construction can, in fact, be done in uniform $\mathsf{TC}^{0}$. Specifically, we observe most of their construction involves sums and products of constantly many $n$-bit integers, which can be done in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. The only involved step is dividing a polynomial of degree (up to) $n$ by a polynomial of degree (up to) $d-1$ and returning the remainder. It turns out that this “polynomial division with remainder” operation can also be performed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (see Corollary 6.5 of Hesse et al., 2002 and an explanation in Section A.3). We thus have the following extension of Mereghetti & Palano’s result:
@@ A.2 Complexity of Matrix Powering
Let $d\in\mathbb{Z}^{+}$ be a fixed constant. Let $\mathbf{M}$ be a $d\times d$ matrix over $n$-bit integers and $z\leq n,z\in\mathbb{Z}^{+}$. Then integer matrix power $\mathbf{M}^{z}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
We extend this to matrix powers over $\mathbb{D}$ rather than integers:
@@
Let $d\in\mathbb{Z}^{+}$ be a fixed constant. Let $\mathbf{M}$ be a $d\times d$ matrix over a datatype $\mathbb{D}$ with entries equivalently representable as $n$-bit integers. Let $z\leq n,z\in\mathbb{Z}^{+}$. If operators $\mathsf{cast}_{\mathbb{D}}$ and $\mathsf{cast}_{\mathbb{R}}$ are in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, then $\mathbb{D}$-matrix power $\mathbf{M}^{z}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
@@
By preconditions of the lemma, we can compute $\mathsf{cast}_{\mathbb{R}}(\mathbf{M})$ in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. Since the entries of $\mathbf{M}$ are equivalent to $n$-bit integers, $\mathsf{cast}_{R}(\mathbf{M})$ can be viewed as a $d\times d$ integer matrix of $n$-bit integers. By Lemma A.6, we can compute $\mathsf{cast}_{R}(\mathbf{M})^{z}$ using an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit. Finally, by a precondition of the lemma, we can cast the result back to $\mathbb{D}$, i.e., compute $\mathsf{cast}_{\mathbb{D}}(\mathsf{cast}_{\mathbb{R}}(\mathbf{M})^{z})$ which equals $\mathbf{M}^{z}$, with an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit. ∎
Lemma A.8 (Float matrix power).
Let $d,c\in\mathbb{Z}^{+}$ be fixed constants. Let $\mathbf{M}$ be a $d\times d$ matrix over $c\log n$ precision floats. Let $z\leq n,z\in\mathbb{Z}^{+}$. Then float matrix power $\mathbf{M}^{z}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
@@
The idea is to convert (by scaling up) $\mathbf{M}$ to another float matrix all whose entries are representable as integers, apply Lemma A.7, reverse the scaling, and cast the result back to $c\log n$ precision floats.
@@
Let $e$ be the smallest exponent across all float entries of $\mathbf{M}$ and $q=\max\{0,-e\}$. Construct a re-scaled float matrix $\tilde{\mathbf{M}}=\mathbf{M}2^{q}$ by adding $q$ to the exponent of every entry of $\mathbf{M}$, using up to $c\log n$ additional bits in the exponent if necessary to keep the computation exact. Note that $e$, $q$, and $\tilde{\mathbf{M}}$ can easily be computed exactly by an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit as they involve fixed-arity arithmetic operations. Further, by construction, $\tilde{\mathbf{M}}$ has non-negative exponents in all its float entries. Thus, every entry of $\tilde{\mathbf{M}}$ represents an integer.
@@
The maximum number representable by each $c\log n$ precision float in $\mathbf{M}$ is upper bounded by $2^{n^{c}}$. Thus, the maximum number representable by each entry of $\tilde{\mathbf{M}}$ is $2^{n^{c}}\times 2^{q}=2^{n^{c}+q}$. Let $m=n^{c}+q$. It follows that each entry $\phi$ of $\tilde{\mathbf{M}}$ can be equivalently represented as an $m$-bit integer. Further, this integer can be computed by left-shifting the mantissa of $\phi$ by a number of bits equal to the value of the exponent of $\phi$ (which is non-negative). Finally, this left-shift, and thus the $\mathsf{cast}_{\mathbb{R}}$ operation over $m$-precision floats, can be easily computed by an $\mathsf{L}$-uniform threshold circuit of size $\mathrm{poly}(m)$. In the other direction, casting from reals to $m$-precision floats can also be easily accomplished by an $\mathsf{L}$-uniform threshold circuit of size $\mathrm{poly}(m)$.
@@
Note that $2^{q}\in[0,n^{c}]$ and hence $m\in[n^{c},2n^{c}]$. In particular, $m\geq n$. Thus $z\leq n$ (a precision) implies $z\leq m$. Observing that $\tilde{\mathbf{M}}$ is a matrix of floats each representable as an $m$-bit integer, we now apply Lemma A.7 with $\mathbb{D}$ being ‘float’ to conclude that float matrix power $\tilde{\mathbf{M}}^{z}$ can be computed by an $\mathsf{L}$-uniform threshold circuit of size $\mathrm{poly}(m)$. Since $m\leq 2n^{c}$, this circuit is also of size $\mathrm{poly}(n)$.
@@
Finally, to compute $\mathbf{M}^{z}$, we first divide each entry of $\tilde{\mathbf{M}}^{z}$ by $2^{qz}$. This can be accomplished by subtracting $qz$ from the exponent of each entry of $\tilde{\mathbf{M}}$; again, we do this computation exactly, using up to $(c+1)\log n$ additional bits in the exponent if necessary. We then cast all entries of the resulting matrix back to $c\log n$ precision floats. All this can be done in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, finishing the proof that $\mathbf{M}^{z}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
A.3 $\mathsf{L}$-Uniformity of Polynomial Division in $\mathsf{TC}^{0}$
@@
Hesse et al. (2002) state that polynomial division is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ in Corollary 6.5. For historical reasons, this claim is preceded by weaker claims in older papers. We briefly clarify this situation to help understand why the stronger claim is valid.
@@
Reif & Tate (1992) establish that polynomial division can be performed in $\mathsf{P}$-uniform $\mathsf{TC}^{0}$, whereas we state our results for $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, which is a smaller class. However, the only issue preventing the polynomial division result from originally going through in the $\mathsf{L}$-uniform case is that, at the time of Reif & Tate’s publication, it was not known whether integer division and iterated integer multiplication are computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. However, Hesse (2001) later proved exactly this. Combining the two results, Theorem 3.2 of Reif & Tate (1992) goes through even with $\mathsf{L}$-uniformity (not just $\mathsf{P}$-uniformity). Its Corollary 3.3 then allows us to conclude that integer polynomial division can be solved by $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuits because the output of integer polynomial division is an analytic function whose Taylor expansion has a finite number of terms (Reif & Tate, 1992).
Appendix B S6 Parameterization
To justify that the S6 architecture used by Mamba is computable in $\mathsf{TC}^{0}$, we justify that $\bar{\mathbf{A}}_{i},\bar{\mathbf{B}}_{i},\mathbf{C}_{i},\mathbf{D}_{i}$ can be computed as a function of $\mathbf{x}_{i}$ in $\mathsf{TC}^{0}$.
We begin by summarizing how exactly is S6 parameterized. S6 first defines continuous-time parameters:
  1. 1.
    $\mathbf{A}$ is a fixed, diagonal matrix that is invertible (each $a_{ii}\neq 0$);
  2. 2.
    $\mathbf{B}_{i}=\pi_{\mathbf{B}}(\mathbf{x}_{i})$ is computed via a projection;
  3. 3.
    $\mathbf{C}_{i}=\pi_{\mathbf{C}}(\mathbf{x}_{i})$ is computed via a projection;
  4. 4.
    $\mathbf{D}_{i}=\mathbf{I}$
Next, we need to discretize the matrices $\mathbf{A}$ and $\mathbf{B}$. S6 does this using an input-dependent discretization factor $\delta_{i}$:
$\delta_{i}=\mathrm{softplus}(\delta+\pi_{\delta}(\mathbf{x}_{i})).$
The discretized matrices are then defined as:
$\bar{\mathbf{A}}_{i}$ $=\exp(\delta_{i}\mathbf{A})$
$\bar{\mathbf{B}}_{i}$ $=(\delta_{i}\mathbf{A})^{-1}\left(\bar{\mathbf{A}}_{i}-\mathbf{I}\right)\delta_{i}\mathbf{B}_{i}.$
It is clear to see that the diagonalizability condition of Theorem 4.4 is satisfied because $\bar{\mathbf{A}}_{i}$ itself is diagonal. Additionally, all the relevant matrices can be computed in $\mathsf{TC}^{0}$.
Proposition B.1.
$\bar{\mathbf{A}}_{i},\bar{\mathbf{B}}_{i},\mathbf{C}_{i},$ and $\mathbf{D}_{i}$ can all be computed as functions of $\mathbf{x}_{i}$ in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
To prove this, observe that $\mathbf{A},\mathbf{B}_{i},\mathbf{C}_{i},\mathbf{D}_{i}$ can all be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ because they are either constants or linear transformations of $\mathbf{x}_{i}$. To justify that $\bar{\mathbf{A}}_{i}$ and $\bar{\mathbf{B}}_{i}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, we just need to justify that we can invert diagonal matrices and compute $\mathrm{softplus}$ and $\exp$ in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Lemma B.2.
Diagonal matrices over log-precision floats can be inverted in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
@@ Appendix B S6 Parameterization
Inverting a diagonal matrix just involves forming the reciprocal along the diagonal. Scalar reciprocals can be approximated to error at most $2^{-n^{c}}$ (for any $c$) in $\mathsf{TC}^{0}$ (Hesse et al., 2002). This means we can compute the reciprocal of a log-precision float (cf. Appendix A) exactly up to log precision. ∎
In Appendix D, we show that we can compute the nonlinearities $\exp$ and $\mathrm{softplus}$ over a bounded domain in $\mathsf{TC}^{0}$.
Appendix C Diagonalizable SSMs
We extend Theorem 4.4 to cover the case when the SSMs transition matrices are simultaneously diagonalizable, rather than just diagonal. This requires us to note that when working with log-precision floating point representations of matrices, a diagonal matrix $\mathbf{A}$ and its diagonalized decomposition $\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$ are numerically substitutable.
See 4.6
Proof.
When the first condition is satisfied, the following equality holds over log-precision floats:
$\prod_{i}\bar{\mathbf{A}}_{i}=\prod_{i}\left(\mathbf{W}\operatorname{diag}(\bar{\mathbf{a}}_{i})\mathbf{W}^{-1}\right).$
By the associativity of $\mathbb{D}$-matrix products, we can remove the parentheses to get
$\prod_{i}\bar{\mathbf{A}}_{i}$ $=\prod_{i}\mathbf{W}\operatorname{diag}(\bar{\mathbf{a}}_{i})\mathbf{W}^{-1}$
$=\mathbf{W}\left[\prod_{i}\operatorname{diag}(\bar{\mathbf{a}}_{i})\right]\mathbf{W}^{-1}.$
Iterated multiplication of diagonal matrices is reducible to several iterated scalar multiplications, which is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Lemma A.5). Then the product of all $\bar{\mathbf{A}}_{i}$ is the product of three $\mathsf{L}$-uniform $\mathsf{TC}^{0}$-computable matrices, so is itself $\mathsf{L}$-uniform $\mathsf{TC}^{0}$-computable. The second condition from Lemma 4.1 is satisfied by assumption. Thus, the convolutional form for $M$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
C.1 Diagonalizable S6
We can define an extension of S6 which satisfies these conditions to show that it is also in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Definition C.1.
Diagonalizable S6 has continuous-time parameters:
  1. 1.
    $\mathbf{A}$ is a fixed matrix diagonalizable as $\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$ that is invertible (each $a_{ii}\neq 0$);
  2. 2.
    $\mathbf{B}_{i}=\pi_{\mathbf{B}}(\mathbf{x}_{i})$ is computed via a projection;
  3. 3.
    $\mathbf{C}_{i}=\pi_{\mathbf{C}}(\mathbf{x}_{i})$ is computed via a projection;
  4. 4.
    $\mathbf{D}=\mathbf{I}$.
As in the standard S6, the discretization of $\mathbf{A}$ and $\mathbf{B}$ is done by an input-dependent discretization factor $\delta_{i}$:
$\delta_{i}=\mathrm{softplus}(\delta+\pi_{\delta}(\mathbf{x}_{i})).$
The discretized matrices are then defined as
$\bar{\mathbf{A}}_{i}$ $=\exp(\delta_{i}\mathbf{A}),$
$\bar{\mathbf{B}}_{i}$ $=(\delta_{i}\mathbf{A})^{-1}(\bar{\mathbf{A}}_{i}-\mathbf{I})\delta_{i}\mathbf{B}_{i}.$
To prove that $\bar{\mathbf{A}}_{i}$ and $\bar{\mathbf{B}}_{i}$ have the necessary properties, we first introduce some lemmas dealing with matrix-valued functions of diagonalizable matrices.
Lemma C.2.
If a matrix $\mathbf{A}$ is diagonalizable, then we can substitute its diagonalized decomposition $\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$ in a computation graph over log-precision floats involving $\mathbf{A}$ without incurring any meaningful error.
Proof.
Let $\mathbf{A}$ be diagonalizable. Then there exists invertible $\mathbf{W}$ and diagonal $\operatorname{diag}(\mathbf{a})$ such that $\mathbf{A}=\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$. Note that the product of a fixed number of matrices is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, and so the first $c\log n$ bits of $\mathbf{A}$ and $\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$ are identical. ∎
Lemma C.3.
Let $\mathbf{A}$ be diagonalizable as $\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$, where $\mathbf{a}\in\mathbb{R}^{d}$. Then $c\cdot\mathbf{A}$ is simultaneously diagonalizable with $\mathbf{A}$ via $c\cdot\mathbf{A}=\mathbf{W}c\cdot\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$.
Proof.
Scalar multiplication commutes around matrix multiplication. ∎
Lemma C.4.
Let $\mathbf{A}$ be diagonalizable as $\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$, where $\mathbf{a}\in\mathbb{R}^{d}$. Then $\exp(\mathbf{A})=\mathbf{W}\exp(\operatorname{diag}(\mathbf{a}))\mathbf{W}^{-1}$.
Proof.
The matrix exponential is defined as a power series, so for diagonalizable $\mathbf{A}$ it follows that
$\exp(\mathbf{A})$ $=\exp(\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1})$
$=\sum_{k=0}^{\infty}\frac{1}{k!}(\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1})^{k}$
$=\sum_{k=0}^{\infty}\frac{1}{k!}\mathbf{W}\operatorname{diag}(\mathbf{a})^{k}\mathbf{W}^{-1}$
$=\mathbf{W}\left(\sum_{k=0}^{\infty}\frac{1}{k!}\operatorname{diag}(\mathbf{a})^{k}\right)\mathbf{W}^{-1}$
$=\mathbf{W}\exp(\operatorname{diag}(\mathbf{a}))\mathbf{W}^{-1}.$
The expressions in Lemma C.4 are equivalent not just over real numbers but also over log-precision floats. This is because we know both expressions can be approximated in $\mathsf{TC}^{0}$ with error at most $2^{-n^{c}}$, which means the $c\log n$ bits of the approximation must be equivalent.
Lemma C.5.
Diagonalizable matrices over log-precision floats can be inverted in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Proof.
Let $\mathbf{A}=\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W^{-1}}$. Then $\mathbf{A}^{-1}=\mathbf{W}^{-1}\operatorname{diag}(\mathbf{a})^{-1}\mathbf{W}$. We are guaranteed that each of these matrices exists, and furthermore by Lemma B.2 we know that $\operatorname{diag}(\mathbf{a})^{-1}$ is computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. Their product, involving a finite number of additions and multiplies, is also computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
Proposition C.6.
$\bar{\mathbf{A}}_{i}$ and $\bar{\mathbf{B}}_{i}$ can be computed as functions of $\mathbf{x}_{i}$ in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Proof.
We first show that $\bar{\mathbf{A}}_{i}$ is $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ computable. By definition,
$\bar{\mathbf{A}}_{i}=\exp(\delta_{i}\mathbf{A}).$
By Corollary D.2, $\delta_{i}$ is computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. The product $\delta_{i}\mathbf{A}$ is simultaneously diagonalizable with $\mathbf{A}$ so
$\bar{\mathbf{A}}_{i}$ $=\exp(\mathbf{W}\delta_{i}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1})$ (Lemma C.3)
$=\mathbf{W}\exp(\operatorname{diag}(\mathbf{a}))\mathbf{W}^{-1}.$ (Lemma C.4)
Since the exponential of scalars is $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ computable by Corollary D.2, then $\bar{\mathbf{A}}_{i}$ is as well.
Turning to $\bar{\mathbf{B}}_{i}$, note that the term $(\delta_{i}\mathbf{A})^{-1}$ is $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ computable by Lemma C.5 since $\delta_{i}\mathbf{A}$ is diagonalizable. Since $\bar{\mathbf{A}}_{i}$ is $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ computable, the difference $\bar{\mathbf{A}}_{i}-\mathbf{I}$ is as well. Then every term in
$\bar{\mathbf{B}}_{i}=(\delta_{i}\mathbf{A})^{-1}(\bar{\mathbf{A}}_{i}-\mathbf{I})\delta_{i}\mathbf{B}_{i}$
is $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ computable, and so their product is as well. ∎
Remark.
Since $\mathbf{C}_{i}$ and $\mathbf{D}_{i}$ are unchanged between the standard and diagonalizable versions of S6, the proofs of their computability as functions of $\mathbf{x}_{i}$ in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ pass through from Appendix B.
Corollary C.7.
There exists an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family that computes Diagonalizable S6’s convolutional form.
Proof.
Note that since $\mathbf{A}=\mathbf{W}\operatorname{diag}(\mathbf{a})\mathbf{W}^{-1}$ is fixed the set of transition matrices $\{\bar{\mathbf{A}}_{i}\}$ is simultaneously diagonalizable via $\mathbf{W}$ for all $i$.
Then Diagonalizable S6 meets the conditions for Theorem 4.6. ∎
Appendix D Nonlinearities in $\mathsf{L}$-Uniform $\mathsf{TC}^{0}$
The parameterization of SSMs (and transformers) involves computing nonlinearities like $\exp$ and $\mathrm{softplus}$. We leverage existing circuit complexity results (Reif & Tate, 1992) to show that, in general, any well-behaved nonlinearity should be computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ when used in conjunction with pre- or post-layer norm.
Lemma D.1 (Adapts Corollary 3.3, Reif & Tate, 1992).
Let $X=(-B,B)$ be a bounded interval. Let $f$ be a function over $X$ with a convergent Taylor series:
$f(x)=\sum_{n=0}^{\infty}\frac{a_{n}}{b_{n}}(x-x_{0})^{n},$
where $a_{n},b_{n}$ are integers with magnitude at most $2^{n^{O(1)}}$ computable in $L$-uniform $\mathsf{TC}^{0}$. Then $f$ can be approximated over $X$ by $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuits to log precision (error at most $2^{-n^{c}}$ for any $c\geq 1$).
Proof.
Reif & Tate (1992) give a proof when $X=(-1,1)$. We generalize to $X=(-B,B)$, assuming w.l.o.g. $B=2^{k}$. The idea is to transform $f$ to have domain $(-1,1)$ via
$g(x)=f(Bx).$
Then, we can apply Corollary 3.3 of Reif & Tate (1992) to approximate $g$ with error at most $2^{-n^{c}}$. Reif & Tate (1992) state their result for $\mathsf{P}$-uniform $\mathsf{TC}^{0}$, but through advances in circuit complexity since the time of publication (Section A.3), their construction naturally applies for $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ as well.
To approximate $f$, compute $z=x/B$, which can be done exactly since $B=2^{k}$. We conclude by computing $g(z)=f(x)$, which, as established, has error at most $2^{-n^{c}}$. ∎
Because of pre- and post-norm layers, the elements of $\mathbf{x}_{i}$ in an SSM will remain in a bounded domain $(-B,B)$. Thus, the following lemma shows we can compute them:
Corollary D.2.
The pointwise nonlinearities $\exp$, $\log$, and $\mathrm{softplus}$ are computable over $(-B,B)$ in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Proof.
By Reif & Tate (1992, Corollary 3.3) know that the Taylor series for $\exp$ and $\log$ is convergent with $a_{n},b_{n}$ computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. Then $\exp$ and $\log$ are themselves computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Since $\mathrm{softplus}(x)=\log\left(1+\exp(x)\right)$ is a fixed composition of $L$-uniform $\mathsf{TC}^{0}$-computable functions, it too is computable in $L$-uniform $\mathsf{TC}^{0}$. ∎